OpenStudy (anguyennn):

How do you solve |x/2-1/5|=1

OpenStudy (anguyennn):

\[\left| \frac{ x }{ 2 }-\frac{ 1 }{ 5 } \right|=1\]

OpenStudy (anguyennn):

hello

OpenStudy (sloppycanada):

First thing first, multiply by the reciprocal.

OpenStudy (anguyennn):

?

OpenStudy (anguyennn):

how do you do that?

OpenStudy (anguyennn):

please help me

OpenStudy (sloppycanada):

\[\frac{ x }{ 2 } x 2 and 1(2)\]

OpenStudy (anguyennn):

??